Given the basis {(1, - 1, 3), (0, 1, - 1),
(0, 3, - 2)} of R3, determine its dual basis.
1
Expert's answer
2020-03-26T07:52:47-0400

Let "e^1=(1,0,0),e^2=(0,1,0),e^3=(0,0,1)" be the standard basis, and "E_1,E_2,E_3" be the its dual basis. Then "\delta^i_j=E_i(e^j)", where "\delta^i_j=\begin{cases}
1,&\text{if $i=j$}\\
0,&\text{if $i\neq j$}
\end{cases}"

Let "f^1=(1,-1,3)=(f^1_1,f^1_2,f^1_3),"

"f^2=(0,1,-1)=(f^2_1,f^2_2,f^2_3), f^3=(0,3,-2)=(f^3_1,f^3_2,f^3_3)" be the given basis, and "F_1=F_1^1E_1+F_1^2E_2+F_1^3E_3,"

"F_2=F_2^1E_1+F_2^2E_2+F_2^3E_3,"

"F_3=F_3^1E_1+F_3^2E_2+F_3^3E_3" be a its dual basis.

We have "E_i(f^j)=E_i(f^j_1e^1+f^j_12e^12+f^j_3e^3)=f^j_i".

Then "\delta_i^j=F_i(f^j)=(F_i^1E_1+F_i^2E_2+F_i^3E_3)(f^j)="

"=F_i^1f^j_1+F_i^2f^j_2+F_i^3f^j_3".

That is "I=\begin{pmatrix}
F_1^1&F_1^2&F_1^3\\
F_2^1&F_2^2&F_2^3\\
F_3^1&F_3^2&F_3^3
\end{pmatrix}\begin{pmatrix}
f_1^1&f_1^2&f_1^3\\
f_2^1&f_2^2&f_2^3\\
f_3^1&f_3^2&f_3^3
\end{pmatrix}", so "\begin{pmatrix}
F_1^1&F_1^2&F_1^3\\
F_2^1&F_2^2&F_2^3\\
F_3^1&F_3^2&F_3^3
\end{pmatrix}=\begin{pmatrix}
f_1^1&f_1^2&f_1^3\\
f_2^1&f_2^2&f_2^3\\
f_3^1&f_3^2&f_3^3
\end{pmatrix}^{-1}"


We obtain "\begin{pmatrix}
F_1^1&F_1^2&F_1^3\\
F_2^1&F_2^2&F_2^3\\
F_3^1&F_3^2&F_3^3
\end{pmatrix}=\begin{pmatrix}
1&0&0\\
-1&1&3\\
3&-1&-2
\end{pmatrix}^{-1}" .

Find "\begin{pmatrix}
1&0&0\\
-1&1&3\\
3&-1&-2
\end{pmatrix}^{-1}".

"\left(\begin{array}{lll|lll}
1&0&0&1&0&0\\
-1&1&3&0&1&0\\
3&-1&-2&0&0&1
\end{array}\right)\to"

"\to\left(\begin{array}{lll|lll}
1&0&0&1&0&0\\
0&1&3&1&1&0\\
0&-1&-2&-3&0&1
\end{array}\right)\to"

"\to\left(\begin{array}{lll|lll}
1&0&0&1&0&0\\
0&1&3&1&1&0\\
0&0&1&-2&1&1
\end{array}\right)\to"

"\to\left(\begin{array}{ccc|ccc}
1&0&0&1&0&0\\
0&1&0&7&-2&-3\\
0&0&1&-2&1&1
\end{array}\right)"

So "\begin{pmatrix}
F_1^1&F_1^2&F_1^3\\
F_2^1&F_2^2&F_2^3\\
F_3^1&F_3^2&F_3^3
\end{pmatrix}=\begin{pmatrix}
1&0&0\\
-1&1&3\\
3&-1&-2
\end{pmatrix}^{-1}=\begin{pmatrix}
1&0&0\\
7&-2&-3\\
-2&1&1
\end{pmatrix}"

That is "F_1=F_1^1E_1+F_1^2E_2+F_1^3E_3=E_1,"

"F_2=F_2^1E_1+F_2^2E_2+F_2^3E_3=7E_1-2E_2-3E_3,"

"F_3=F_3^1E_1+F_3^2E_2+F_3^3E_3=-2E_1+E_2+E_3"

Answer: "E_1, 7E_1-2E_2-3E_3, -2E_1+E_2+E_3"